### 01 Apr 2026 v1

# Verifying Well-Posedness of Linear PDEs using Convex Optimization

Abstract
Ensuring that a PDE model is well-posed is a necessary precursor to any form of analysis, control, or numerical simulation. Although the Lumer-Phillips theorem provides necessary and sufficient conditions for well-posedness of dissipative PDEs, these conditions must hold only on the domain of the PDE – a proper subspace of $L_{2}$ – which can make them difficult to verify in practice. In this paper, we show how the Lumer–Phillips conditions for PDEs can be tested more conveniently using the equivalent Partial Integral Equation (PIE) representation. This representation introduces a fundamental state in the Hilbert space $L_{2}$ and provides a bijection between this state space and the PDE domain. Using this bijection, we reformulate the Lumer-Phillips conditions as operator inequalities on $L_{2}$. We show how these inequalities can be tested using convex optimization methods, establishing a least upper bound on the exponential growth rate of solutions. We demonstrate the effectiveness of the proposed approach by verifying well-posedness for several classical examples of parabolic and hyperbolic PDEs.
Acknowledgement: This work was supported by National Science Foundation grants 2337751 and 2429973.
I INTRODUCTION
Well-Posedness is a fundamental property of any dynamic model, ensuring existence and regularity of solutions. For ordinary differential equations, the problem is relatively simple. Any linear ordinary differential equation is well-posed and even for nonlinear ordinary differential equations, existence of a non-increasing Lyapunov function is sufficient to establish this property. Well-posedness of Partial Differential Equations (PDEs), however, is more nuanced. Specifically, uniqueness of solutions necessitates the imposition of boundary conditions (and the associated spatial Sobolev regularity required to ensure existence of such boundary values). The well-posedness condition, then, requires that the boundary conditions (BCs) are tight enough to avoid multiple solutions, while not being so tight as to preclude any solution. This process of ensuring that the BCs are compatible with the PDE is often a difficult time-consuming process, requiring careful ad hoc analysis and integration of domain (BCs) and generator (PDE). In this paper, we propose a convex optimization algorithm for testing well-posedness which applies to a large class of linear PDEs and BCs, while avoiding the need for extensive ad hoc analysis.
Let us consider a linear PDE of the form
$\dot{\mathbf{u}}(t)=A\mathbf{u}(t),\quad t\geq 0,\qquad\mathbf{u}(0)=\mathbf{x},$ (1)
where $\mathbf{u}(t)\in D(A)$ denotes the state of the system, defined on a Sobolev subspace $D(A)\subseteq H$ of a Hilbert space $H$, and where $A:D(A)\to H$ is a linear differential operator. Many physical phenomena can be modeled by dynamical systems of this form. Such models provide simplified descriptions of complex processes (e.g. physical, chemical, biological, etc.), enabling analysis of system properties such as stability and facilitating the synthesis of controllers and observers.
A PDE model of the form in Eqn. (1) is well posed if it admits a unique solution that depends continuously on the initial state $\mathbf{x}$ [1]. A common framework for formulating well-posedness of PDEs is the theory of $C_{0}$-semigroups [2]. These semigroups generalize the matrix exponential to infinite-dimensional settings, allowing solutions of (1) to be characterized by the $C_{0}$-semigroup generated by $A$. Within this framework, the question of well-posedness is whether the operator $A$ generates a $C_{0}$-semigroup, and substantial research has been devoted to establishing generation theorems which provide conditions under which such a semigroup exists.
While necessary and sufficient conditions for generation of $C_{0}$-semigroups have been established in the form of the Hille–Yosida theorem [^1] Also known as the Feller–Miyadera–Phillips theorem, for the more general formulation derived in [3, 4, 5]. [6, 7], the formulation of these conditions is abstract and verification is challenging. Consequently, most well-posedness results use the stronger conditions proposed in the Lumer–Phillips theorem [8]. These conditions are more closely related to the well-posedness test for nonlinear ODEs based on existence of non-increasing Lyapunov function. Specifically, the Lumer-Phillips theorem states that $A$ generates a contraction semigroup – implying that some norm of solutions in some Hilbert space does not increase with time – if and only if $A$ satisfies $\left\langle\mathbf{x},A\mathbf{x}\right\rangle_{H}\leq 0$ for all $\mathbf{x}\in D(A)$, and $\lambda I-A$ is surjective for some $\lambda>0$. Modifications of the Lumer–Phillips theorem have also been proposed for other classes of systems, including time-varying PDEs [9, 10] and partial differential–algebraic equations [11, 12].
Despite being widely used to verify well-posedness of PDEs, application of the Lumer–Phillips theorem is typically tailored to the specific system under consideration. Development of a universal framework for testing well-posedness is complicated by the lack of a universal parameterization of operators $A$ defining PDEs on different domains $D(A)$. As a result, substantial analysis is often required to test the conditions of the Lumer–Phillips theorem. For instance, dissipativity is commonly verified using techniques such as integration by parts and the Cauchy–Schwarz inequality to establish that $\left\langle\mathbf{x},A\mathbf{x}\right\rangle_{H}\leq 0$ for all $\mathbf{x}\in D(A)$. Similarly, surjectivity of $\lambda I-A$ is frequently tested by explicitly inverting the operator $\lambda I-A$ or $A$, which becomes highly nontrivial for more complex systems.
To develop a more general, computational approach to well-posedness analysis of PDEs, we propose to use the Partial Integral Equation (PIE) representation. This PIE representation provides an equivalent state-space formulation of PDEs by establishing a bijection between the PDE domain and the Hilbert space $H=L_{2}$. Under suitable rank conditions on the boundary conditions defining $D(A)$ (see [13], Sec. XIV.3, Thm. 3.1), such a bijection exists and belongs to a class of Partial Integral (PI) operators of the form
$\mathcal{T}\mathbf{v}(s):=\int_{0}^{s}\boldsymbol{T}_{1}(s,\theta)\mathbf{v}(\theta)d\theta+\int_{s}^{1}\boldsymbol{T}_{2}(s,\theta)\mathbf{v}(\theta)d\theta,$
where the kernels $\boldsymbol{T}_{1}$ and $\boldsymbol{T}_{2}$ are polynomial and can be computed explicitly from the boundary conditions defining $D(A)$. Using this operator, $\mathcal{T}:L_{2}\to D(A)$ is bijective, yielding an equivalent representation of the PDE in (1) as
$\partial_{t}\mathcal{T}\mathbf{v}(t)=\mathcal{A}\mathbf{v}(t),\quad t\geq 0,\qquad\mathbf{v}(0)=\mathcal{T}^{-1}\mathbf{x},$
where $\mathcal{A}:=A\circ\mathcal{T}^{-1}$ is also a PI operator. It has been shown that a broad class of linear ODE–PDE systems can be represented equivalently as PIEs of this form [14, 15], where $\mathbf{u}(t)$ satisfies the PDE if and only if $\mathbf{u}(t)=\mathcal{T}\mathbf{v}(t)$ for some solution $\mathbf{v}(t)$ of the PIE.
Within the PIE framework, various analysis and synthesis problems – including stability analysis as well as observer and controller design – can be formulated as convex optimization problems over PI operator variables, known as Linear PI Inequalities (LPIs). Since PI operators admit polynomial parameterizations, these LPIs can be solved using semidefinite programming methods through the PIETOOLS software suite [16].
Despite the growing use of the PIE representation for analysis and control of PDE systems, a corresponding generation theorem has not yet been formulated in the PIE framework. The presence of the operator $\mathcal{T}$ on the left-hand side of the PIE dynamics prevents direct application of classical results such as the Lumer–Phillips theorem, raising the question of how well-posedness conditions should be formulated in terms of the operators $\{\mathcal{T},\mathcal{A}\}$. Although related generation results exist for descriptor systems of the form $\frac{d}{dt}E\mathbf{x}(t)=A\mathbf{x}(t)$ [11, 12], these results still assume $A$ to be defined on a domain $D(A)$, leading to conditions that are not significantly easier to verify than the classical generator conditions. This leaves the challenge of how to exploit the bijective nature of $\mathcal{T}$ and the fact that $\mathcal{A}\in\mathcal{L}(H)$ to establish a simplified generation theorem, of which the conditions can be tested numerically.
In this paper, we address this challenge by deriving well-posedness conditions directly in the PIE representation. First, in Section III, we parameterize a class of 1D, linear PDEs admitting an equivalent PIE representation, and formulate the problem of verifying well-posedness on $L_{2}$ in terms of the associated PIE. In Section IV, we then formulate a Lumer–Phillips theorem for PIEs, obtaining necessary and sufficient conditions for the PIE to generate a quasicontraction semigroup. Finally, in Section V, we use this result to show how well-posedness of a PDE can be tested in the PIE representation by solving an LPI, further establishing a bound on the exponential growth rate of solutions. We apply this test to several classical PDE examples in Section VI, using the PIETOOLS software to verify well-posedness and compute tight bounds on the growth rate of solutions.
II Preliminaries
II-A Notation
Define $\mathbb{R}_{+}:=[0,\infty)$. For $d,n\in\mathbb{N}$, denote by $L_{2}^{n}[a,b]$ the Hilbert space of $\mathbb{R}^{n}$-valued square-integrable functions on $[a,b]$, with the standard inner product $\left\langle\cdot,\cdot\right\rangle_{L_{2}}$, and where we omit the domain when clear from context. For $\mathcal{P}\in L_{2}^{n}[a,b]\to L_{2}^{n}[a,b]$ symmetric positive definite, define the weighted inner product $\left\langle\mathbf{x},\mathbf{y}\right\rangle_{\mathcal{P}}:=\left\langle\mathbf{x},\mathcal{P}\mathbf{y}\right\rangle_{L_{2}}$ for $\mathbf{x},\mathbf{y}\in L_{2}^{n}$. Define the Sobolev subspace
$W_{2}^{d,n}[0,1]:=\{\mathbf{u}\in L_{2}^{n}\mid\partial_{s}^{k}\mathbf{u}\in L_{2}^{n}[0,1],~\forall k\in\{1,\ldots,d\}\}.$
For a Hilbert space $H$, denote by $\mathcal{L}(H)$ the space of bounded linear operators on $H$ (with operator norm induced by $\left\lVert{\cdot}\right\rVert_{H}$), and let $I\in\mathcal{L}(H)$ denote the identity operator. For $\mathcal{P}\in\mathcal{L}(H)$, denote by $\textnormal{ran}(\mathcal{P})$ and $\textnormal{ker}(\mathcal{P})$ the range and nullspace of $\mathcal{P}$, respectively.
II-B Strongly Continuous Semigroups
Well-posedness of solutions to linear PDEs as in (1) will be verified using the theory of $C_{0}$-semigroups. For this, we recall the following definitions.
Definition 1 (Infinitesimal Generator of $C_{0}$-Semigroup)
For a Hilbert space $H$, a $\boldsymbol{C_{0}}$-semigroup (or strongly continuous semigroup) on $H$ is a family $(S(t))_{t\geq 0}$ of operators in $\mathcal{L}(H)$ with the following properties:
  1. 1.
    $S(0)=I$;
  2. 2.
    $S(t+s)=S(t)S(s)$ for all $t,s\geq 0$;
  3. 3.
    $\lim_{t\to 0^{+}}S(t)\mathbf{x}=\mathbf{x}$ for all $\mathbf{x}\in H$.
For a subspace $D\subseteq H$, an operator $A:D\to H$ is the infinitesimal generator of $(S(t))_{t\geq 0}$ if for all $\mathbf{x}\in D$,
$A\mathbf{x}=\lim_{t\to 0^{+}}\frac{1}{t}[S(t)-I]\mathbf{x}.$
In this case, we define $e^{tA}:=S(t)$ for $t\geq 0$.
It is well-established that if $A$ is the infinitesimal generator of a $C_{0}$-semigroup, then solutions to the PDE in (1) defined by AA take the form $\mathbf{u}(t)=e^{tA}\mathbf{x}$, for initial state $\mathbf{x}\in D$ [17]. Accordingly, in the following section, well-posedness of the PDE will be defined in terms of generation of a $C_{0}$-semigroup.
III Well-Posedness in the PIE Representation
The main contribution of this paper is a convex optimization program for verifying well-posedness of a class of linear PDEs. This optimization program will be established using the equivalent Partial Integral Equation (PIE) representation of the PDE, modeling the highest-order spatial derivative of the PDE state. In this section, we recall how this PIE representation may be constructed for a broad class of linear, 1D PDEs, formulating the problem of well-posedness analysis in terms of the associated PIE representation.
III-A Bijection between PDE Domain and $L_{2}$
We begin by defining a class of domains compatible with the PIE representation.
Definition 2
For $n,d\in\mathbb{N}$, $a,b]\subseteq\mathbb{R}$, and $B,C\in\mathbb{R}^{nd\times nd}$, define the associated domain $D\subseteq W_{2}^{d,n}[a,b]$ as
$D:=\biggl\{\mathbf{u}\in W_{2}^{d,n}[a,b]\,\biggl|\ \sum_{j=1}^{d}B_{i,j}\partial_{s}^{j}\mathbf{u}(a)+C_{i,j}\partial_{s}^{j}\mathbf{u}(b)=0\biggr\},$
where $B_{i,j},C_{i,j}\in\mathbb{R}^{n\times n}$ are such that $B=[B_{i,j}]$ and $C=[C_{i,j}]$. We say that $D$ is PIE compatible if the matrix $K:=B+C\boldsymbol{Q}(b-a)$ is invertible, where
$\boldsymbol{Q}(z):={\small\begin{bmatrix}I_{n}&zI_{n}&\cdots&\frac{z^{d-1}}{(d-1)!}I_{n}\\ 0_{n}&I_{n}&\ddots&\frac{z^{d-2}}{(d-2)!}I_{n}\\ \vdots&\ddots&\ddots&\vdots\\ 0_{n}&0_{n}&\cdots&I_{n}\end{bmatrix}}$.
Many common boundary conditions (e.g., Dirichlet, Robin, mixed) yield PIE compatible domains. The following result, a corollary of Thm. 3.1 in Sec. XIV.3 of [13], shows that for any such PIE compatible domain, the operator $\partial_{s}^{d}:D\to L_{2}^{n}[a,b]$ is invertible.
Corollary 3
If $D\subseteq W_{2}^{d,n}[a,b]$ is PIE compatible, then $\partial_{s}^{d}:D\to L_{2}^{n}[a,b]$ is invertible, and $\mathcal{T}:=(\partial_{s}^{d})^{-1}:L_{2}^{n}[a,b]\to D$ takes the form $(\mathcal{T}\mathbf{v})(s):=\int_{a}^{b}\boldsymbol{G}(s,\theta)\mathbf{v}(\theta)\,d\theta,$ where
$\boldsymbol{G}(s,\theta):=\begin{cases}E_{1}^{T}\boldsymbol{Q}(s-a)(I-K^{-1}C)\boldsymbol{Q}(b-\theta)E_{d},&\theta\leq s,\\ -E_{1}^{T}\boldsymbol{Q}(s-a)K^{-1}C\boldsymbol{Q}(b-\theta)E_{d},&\theta>s,\end{cases}$
for $\boldsymbol{Q}(s)$ and $K$ as in Defn. 2, and $E_{1}=\begin{bmatrix}I_{n}&0_{n}&\cdots&0_{n}\end{bmatrix}^{T}$ and $E_{d}=\begin{bmatrix}0_{n}&\cdots&0_{n}&I_{n}\end{bmatrix}^{T}$.
Cor. 3 shows that if a PDE domain $D\subseteq L_{2}^{n}$ is PIE compatible as in Defn. 2, then there exists a unique solution to the boundary value problem
$\partial_{s}^{n}\mathbf{u}=\mathbf{v},\qquad\mathbf{u}\in D,$ (2)
for all $\mathbf{v}\in D$. Moreover, this solution can be explicitly constructed as $\mathbf{u}=\mathcal{T}\mathbf{v}$, where $\mathcal{T}$ is a bounded, linear operator on $L_{2}^{n}$. In this manner, for a linear 1D PDE, PIE compatibility of the PDE domain already establishes well-posedness of the underlying boundary value problem in (2). This will greatly simplify well-posedness analysis of the full PDE, by using the operator $\mathcal{T}$ to establish a state space representation of the PDE in terms of $\mathbf{v}(t)\in L_{2}^{n}$—as a PIE.
III-B The PIE Representation of Linear 1D PDEs
Given the class of PIE compatible domains from the previous subsection, we now consider the challenge of verifying well-posedness of a linear PDE defined on such a domain. In particular, we consider the following abstract Cauchy problem on $D\subseteq L_{2}^{n}[a,b]$,
$\dot{\mathbf{u}}(t)=A\mathbf{u}(t),\quad t\geq 0,\qquad\mathbf{u}(0)=\mathbf{x},$ (3)
where $\mathbf{x}\in D$ is the initial state, and where we define a class of linear operators
$(A\mathbf{u})(s):=\sum_{k=0}^{d}\boldsymbol{A}_{k}(s)\partial_{s}^{k}\mathbf{u}(s),\quad\mathbf{u}\in D,$ (4)
parameterized by $\boldsymbol{A}_{k}\in\mathbb{R}^{n\times n}[s]$ for $k\in\{0,\ldots,d\}$. Note that this parameterization supports a broad class of linear, 1D PDEs, several examples of which are provided in Sec. VI. Given this class of PDEs, we will consider the following definition of well-posedness throughout the paper.
Definition 4 (Well-Posed PDE)
For a given subspace $D\subseteq L_{2}^{n}[a,b]$ and $A:D\to L_{2}^{n}[a,b]$, we say that the PDE defined by $A$ is well-posed if $A$ is the infinitesimal generator of a $C_{0}$-semigroup on $L_{2}^{n}[a,b]$.
Although not immediately apparent from the definition, well-posedness of a given PDE depends not only on the operator $A$, but also on the domain $D$ and the underlying Hilbert space $L_{2}^{n}[a,b]$ on which the system is defined. This distinction is particularly important for PDE models that admit multiple representations in terms of different operators $A$ and domains DD—such as the wave equation in Subsection VI-C—for which well-posedness of one representation does not necessarily imply well-posedness of another. The proposed parameterizations of $D$ (Defn. 2) and $A$ (Eqn. 4) will allow for verification of well-posedness for different such representations. However, well-posedness will only be verified on the Hilbert space $L_{2}^{n}[a,b]$, with extensions to alternative Hilbert spaces left for future study.
To verify well-posedness of PDEs as in (3), recall that if the PDE domain $D$ is PIE-compatible, then any element of $D$ may be uniquely identified with an element of $L_{2}^{n}$ using the bijection $\mathcal{T}:=(\partial_{s}^{d})^{-1}:L_{2}^{n}\to D$. Introducing the state transformation $\mathbf{u}(t)=\mathcal{T}\mathbf{v}(t)$, then, the PDE in (3) can be equivalently represented as a PIE, taking the form
$\partial_{t}\mathcal{T}\mathbf{v}(t)=\mathcal{A}\mathbf{v}(t),\quad t\geq 0,\qquad\mathbf{v}(0)=\mathcal{T}^{-1}\mathbf{x},$ (5)
where $\mathcal{A}\in\mathcal{L}(L_{2}^{n}[a,b])$ is defined by
$\mathcal{A}:=A\circ\mathcal{T}:=\sum_{k=0}^{d}\textnormal{M}_{\boldsymbol{A}_{k}}\circ\partial_{s}^{k}\circ\mathcal{T},$
for $\textnormal{M}_{\boldsymbol{A}}$ the multiplier operator defined by $\boldsymbol{A}$. It is not difficult to see that for any solution to the PDE, there exists an associated solution to the PIE, and vice versa, establishing equivalence of the two representations [14]. As such, a notion of well-posedness of the PIE can also be defined in terms of the associated PDE as follows.
Definition 5 (Well-Posed PIE)
For $\mathcal{T}:L_{2}^{n}[a,b]\to D$ and $\mathcal{A}\in\mathcal{L}(L_{2}^{n}[a,b])$ bounded linear operators, with $\mathcal{T}$ bijective onto $D$, we say that the PIE defined by $\{\mathcal{T},\mathcal{A}\}$ is well-posed if $A:=\mathcal{A}\circ\mathcal{T}^{-1}:D\to L_{2}^{n}$ is the infinitesimal generator of a $C_{0}$-semigroup. We refer to the $C_{0}$-semigroup generated by $\{\mathcal{T},\mathcal{A}\}$ as that generated by $A$.
By this definition, well-posedness of a given PIE is equivalent to well-posedness of its associated PDE. The key advantage of the PIE formulation, however, is that both $\mathcal{T}$ and $\mathcal{A}$ are bounded operators on $L_{2}$, and the boundary conditions and regularity constraints are embedded in $\mathcal{T}$. This will allow conditions for well-posedness of the PIE to be formulated in terms of operator inequalities and tested using convex optimization, as we show in the following sections.
IV A Lumer–Phillips Theorem for PIEs
To establish conditions for well-posedness in the PIE representation, in this section, we first propose a PIE-based formulation of the classical Lumer–Phillips generation theorem. This result may be used to verify generation of contraction and quasicontraction semigroups, defined as follows.
Definition 6 (Contraction Semigroup)
For a Hilbert space $H$, a $C_{0}$-semigroup $(S(t))_{t\geq 0}$ on $H$ is a quasocontraction semigroup with rate $\omega\in\mathbb{R}$ if $\left\lVert{S(t)}\right\rVert_{H}\leq e^{\omega t}$ for all $t\geq 0$. A contraction semigroup on $H$ is a quasicontraction semigroup with rate $\omega=0$.
Generators of contraction semigroups correspond to PDEs for which the origin is Lyapunov stable, so that the functional $V(\mathbf{u}):=\left\lVert{\mathbf{u}}\right\rVert_{H}^{2}$ is non-increasing along solutions. This requires the generator $A:D\to H$ to be dissipative, i.e., $\left\langle\mathbf{u},A\mathbf{u}\right\rangle\leq 0$ for all $\mathbf{u}\in D$—a condition that is also sufficient for well-posedness of (nonlinear) ODEs. For PDEs, however, an additional condition is required to ensure that $e^{tA}$ is well-defined, as provided by the Lumer–Phillips theorem (Thm. 3.4.2 of [18]).
Theorem 7 (Lumer–Phillips Theorem)
For a given Hilbert space $H$ and subspace $D\subseteq H$, let $A:D\to H$ be a linear operator. Then $A$ is the infinitesimal generator of a contraction semigroup on $H$ if and only if $\left\langle\mathbf{u},A\mathbf{u}\right\rangle_{H}\leq 0$ for all $\mathbf{u}\in D$ and the operator $\lambda I-A:D\to H$ is surjective for any (or equivalently all) $\lambda>0$.
Although the Lumer–Phillips theorem only characterizes generators of contraction semigroups, it is well-known that $A$ is the generator of a quasicontraction semigroup with rate $\omega$ if and only if $A-\omega I$ is the generator of a contraction semigroup [17] – making the Lumer–Phillips a powerful tool for verifying generation of quasicontraction semigroups. However, while various techniques have been developed to verify the dissipativity condition of the Lumer–Phillips theorem, testing surjectivity of $\lambda I-A$ for $\lambda>0$ is typically more challenging. In the following subsections, we reformulate both of these conditions in terms of the operators $\{\mathcal{T},\mathcal{A}\}$ defining the associated PIE, using the fact that $\mathcal{T}:L_{2}^{n}\to D$ is bijective to show that both conditions can be more conveniently tested in the PIE representation. In the subsequent subsection, we then use these conditions to established a PIE-based Lumer–Phillips theorem, providing necessary and sufficient conditions for $\{\mathcal{T},\mathcal{A}\}$ to generate a quasicontraction semigroup.
Throughout this section, we will consider a general Hilbert space $H$. In the following section, this freedom in the choice of $H$ will be exploited to verify generation of a more general $C_{0}$-semigroup on $L_{2}$, by choosing $H=(L_{2},\left\langle\cdot,\cdot\right\rangle_{\mathcal{P}})$ for a coercive and bounded operator $\mathcal{P}$.
IV-A Dissipativity in the PIE Representation
Consider first the dissipativity condition, $\left\langle\mathbf{u},A\mathbf{u}\right\rangle_{H}\leq 0$ for all $\mathbf{u}\in D$. This condition is equivalent to Lyapunov stability of the origin for the PDE, conditions for which have already been established in the associated PIE representation, and may be formulated as in the following lemma.
Lemma 8
For a Hilbert space $H$, let $\mathcal{T}:H\to D$ and $\mathcal{A}:H\to H$ be bounded linear operators, with $\mathcal{T}$ bijective onto $D$. Define $A:=\mathcal{A}\circ\mathcal{T}^{-1}:D\to H$. Then $\left\langle\mathbf{u},A\mathbf{u}\right\rangle_{H}\leq 0$ for all $\mathbf{u}\in D$ if and only if $\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{H}\leq 0$ for all $\mathbf{v}\in H$.
Proof:
For sufficiency, suppose $\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{H}\leq 0$ for all $\mathbf{v}\in H$. Then, for any $\mathbf{u}\in D$, letting $\mathbf{v}:=\mathcal{T}^{-1}\mathbf{u}\in H$,
$\left\langle\mathbf{u},A\mathbf{u}\right\rangle_{H}=\left\langle\mathcal{T}\mathbf{v},(\mathcal{A}\circ\mathcal{T}^{-1})\mathcal{T}\mathbf{v}\right\rangle_{H}=\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{H}\leq 0.$
Similarly, for necessity, suppose $\left\langle\mathbf{u},A\mathbf{u}\right\rangle_{H}\leq 0$ for all $\mathbf{u}\in D$. Then, for any $\mathbf{v}\in H$, letting $\mathbf{u}:=\mathcal{T}\mathbf{v}\in D$,
$\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{H}=\left\langle\mathbf{u},(A\circ\mathcal{T})\mathcal{T}^{-1}\mathbf{u}\right\rangle_{H}=\left\langle\mathbf{u},A\mathbf{u}\right\rangle_{H}\leq 0.$
By Lem. 8, dissipativity of $A=\mathcal{A}\circ\mathcal{T}^{-1}$ can be verified by testing $\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{H}\leq 0$ for all $\mathbf{v}\in H$. Here, while AA is defined only on $D$, the operators $\mathcal{T}$ and $\mathcal{A}$ act on all of $H$. Using the inner product $\left\langle\cdot,\cdot\right\rangle_{H}:=\left\langle\cdot,\cdot\right\rangle_{\mathcal{P}}$ for some operator $\mathcal{P}$, this will allow dissipativity to be tested as an operator inequality on $\mathcal{T},\mathcal{A},\mathcal{P}$, as shown in Subsec. V.
IV-B Surjectivity in the PIE Representation
Consider now the second condition of the Lumer–Phillips theorem: surjectivity of $\lambda I-A:D\to H$ for some $\lambda>0$. Although verifying surjectivity of $\lambda I-A$ for general differential operators $A:W_{2}^{d,n}[a,b]\to L_{2}^{n}[a,b]$ is often nontrivial, Cor. 3 shows that for $D\subseteq W_{2}^{d,n}$ parameterized as in Defn. 2, surjectivity of $\partial_{s}^{d}:D\to L_{2}^{n}$ is implied by PIE compatibility of $D$, with an explicit inverse given by $\mathcal{T}$. Using this fact, sujrectivity of $\lambda I-A=(\lambda\mathcal{T}-\mathcal{A})\mathcal{T}^{-1}:D\to L_{2}^{n}$ is then equivalent to surjectivity of $(\lambda\mathcal{T}-\mathcal{A}):L_{2}^{n}\to L_{2}^{n}$, for any $\lambda>0$. Unlike $\lambda I-A$, however, $\lambda\mathcal{T}-\mathcal{A}$ is a bounded operator, allowing surjectivity to be analysed using the open mapping theorem. In particular, the following lemma (a reformulation of Thm. 4.13 in [19]) shows that this property is equivalent to $(\lambda\mathcal{T}-\mathcal{A})^{*}$ being bounded below.
Lemma 9
Let $\mathcal{B}\in\mathcal{L}(H)$ for a Hilbert space $H$. Then $\mathcal{B}$ is surjective if and only if there exists $\epsilon>0$ such that
$\left\lVert{\mathcal{B}^{*}\mathbf{x}}\right\rVert_{H}\geq\epsilon\left\lVert{\mathbf{x}}\right\rVert_{H},\qquad\forall\mathbf{x}\in H.$
Proof:
For sufficiency, suppose there exists $\epsilon>0$ such that $\left\lVert{\mathcal{B}^{*}\mathbf{x}}\right\rVert_{H}\geq\epsilon\left\lVert{\mathbf{x}}\right\rVert_{H}$ for all $\mathbf{x}\in H$. Then $\left\lVert{\mathcal{B}^{*}\mathbf{x}}\right\rVert_{H}=0$ implies $\mathbf{x}=0$, whence $\textnormal{ker}(\mathcal{B}^{*})=\{0\}$. By the closed range theorem, in order to prove $\textnormal{ran}(\mathcal{B})=\textnormal{ker}(\mathcal{B}^{*})^{\perp}=H$, it suffices to show that $\textnormal{ran}(\mathcal{B}^{*})$ is closed.
To show that $\textnormal{ran}(\mathcal{B}^{*})$ is closed, let $(\mathcal{B}^{*}\mathbf{x}_{n})$ be a convergent sequence in $\textnormal{ran}(\mathcal{B}^{*})$ with limit $\mathbf{y}$. Then, for all $m,n\in\mathbb{N}$,
$\left\lVert{\mathbf{x}_{n}-\mathbf{x}_{m}}\right\rVert_{H}\leq\frac{1}{\epsilon}\left\lVert{\mathcal{B}^{*}\mathbf{x}_{n}-\mathcal{B}^{*}\mathbf{x}_{m}}\right\rVert_{H}.$
Since $(\mathcal{B}\mathbf{x}_{n})$ is Cauchy, also $(\mathbf{x}_{n})$ is Cauchy, and therefore converges to some $\mathbf{x}\in H$. By continuity of $\mathcal{B}^{*}$, this implies $\mathcal{B}^{*}\mathbf{x}=\mathbf{y}$, whence $\textnormal{ran}(\mathcal{B}^{*})$ is closed and $\textnormal{ran}(\mathcal{B})=H$.
For necessity, suppose $\textnormal{ran}(\mathcal{B})=H$. By the open mapping theorem, $\mathcal{B}$ is open. Let $U$ denote the open unit ball in $H$. Then there exists $\epsilon>0$ such that $\epsilon U\subseteq\mathcal{B}(U)$. It follows that, for all $\mathbf{x}\in H$,
$\left\lVert{\mathcal{B}^{*}\mathbf{x}}\right\rVert_{H}$ $=\sup_{\mathbf{y}\in U}|\left\langle\mathcal{B}^{*}\mathbf{x},\mathbf{y}\right\rangle_{H}|$
$=\sup_{\mathbf{y}\in U}|\left\langle\mathbf{x},\mathcal{B}\mathbf{y}\right\rangle_{H}|\geq\sup_{\mathbf{z}\in\epsilon U}|\left\langle\mathbf{x},\mathbf{z}\right\rangle_{H}|=\epsilon\left\lVert{\mathbf{x}}\right\rVert_{H}.$
By Lemma 9, surjectivity of $\lambda\mathcal{T}-\mathcal{A}:L_{2}^{n}\to L_{2}^{n}$ can be tested by verifying that $\lambda\mathcal{T}^{*}-\mathcal{A}^{*}$ is bounded below. As for the dissipativity constraint, this condition can be posed as on operator inequality and tested using semidefinite programming, as we show in Sec. V.
IV-C A PIE-Based Generation Theorem for Quasicontraction Semigroups
Having shown how both conditions of the Lumer–Phillips theorem can be tested in the PIE representation, we now combine these results to obtain a PIE-based formulation of the Lumer–Phillips theorem. Here, we remark that if $A:D\to H$ generates a $C_{0}$-semigroup on $H$, then for any $\omega\in\mathbb{R}$, $B:=A+\omega I:D\to H$ also generates a $C_{0}$-semigroup, satisfying $e^{tB}=e^{\omega t}e^{tA}$ for all $t\geq 0$ [17]. Using this observation, and applying Lem. 8 and Lem. 9, we obtain the following necessary and sufficient conditions for $\{\mathcal{T},\mathcal{A}\}$ to generate a quasicontraction semigroup.
Theorem 10
For a given Hilbert space $H$, let $\mathcal{T}:H\to D$ and $\mathcal{A}\in\mathcal{L}(H)$ be bounded linear operators, with $\mathcal{T}$ bijective onto $D$. Define $A:=\mathcal{A}\circ\mathcal{T}^{-1}:D\to H$. Then for any $\omega\in\mathbb{R}$, $A$ is the infinitesimal generator of a quasicontraction semigroup on $H$ with rate $\omega$ if and only if
  1. 1.
    $\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{H}\leq\omega\left\langle\mathcal{T}\mathbf{v},\mathcal{T}\mathbf{v}\right\rangle_{H}$ for all $\mathbf{v}\in H$;
  2. 2.
    there exist $\epsilon>0$ and $\lambda>\omega$ such that $\left\lVert{(\lambda\mathcal{T}^{*}-\mathcal{A}^{*})\mathbf{v}}\right\rVert_{H}\geq\epsilon\left\lVert{\mathbf{v}}\right\rVert_{H}$ for all $\mathbf{v}\in H$.
Proof:
For sufficiency, suppose both conditions hold. Define $A_{\omega}:=A-\omega I:D\to H$ and $\mathcal{A}_{\omega}:=A_{\omega}\circ\mathcal{T}=\mathcal{A}-\omega\mathcal{T}\in\mathcal{L}(H)$. Then for all $\mathbf{v}\in H$,
$\left\langle\mathcal{T}\mathbf{v},\mathcal{A}_{\omega}\mathbf{v}\right\rangle_{H}=\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{H}-\omega\left\langle\mathcal{T}\mathbf{v},\mathcal{T}\mathbf{v}\right\rangle_{H}\leq 0.$
By Lem. 8, it follows that $\left\langle\mathbf{u},A_{\omega}\mathbf{u}\right\rangle_{H}\leq 0$ for all $\mathbf{u}\in D$. In addition, if $\lambda>\omega$ satisfies $\left\lVert{(\lambda\mathcal{T}^{*}-\mathcal{A}^{*})\mathbf{v}}\right\rVert_{H}\geq\epsilon\left\lVert{\mathbf{v}}\right\rVert_{H}$, then for $\lambda_{\omega}:=\lambda-\omega>0$
$\left\lVert{(\lambda_{\omega}\mathcal{T}^{*}-\mathcal{A}_{\omega}^{*})\mathbf{v}}\right\rVert_{H}=\left\lVert{(\lambda\mathcal{T}^{*}-\mathcal{A}^{*})\mathbf{v}}\right\rVert_{H}\geq\epsilon\left\lVert{\mathbf{v}}\right\rVert_{H}.$
By Lem. 9, it follows that $(\lambda\mathcal{T}-\mathcal{A}_{\omega})\in\mathcal{L}(H)$ is surjective. Since $\mathcal{T}:H\to D$ is bijective, this implies $(\lambda I-A_{\omega}):=(\lambda\mathcal{T}-A_{\omega})\circ\mathcal{T}^{-1}:D\to H$ is surjective. By the Lumer–Phillips theorem (Thm. 7), it follows that $A_{\omega}=\mathcal{A}_{\omega}\circ\mathcal{T}^{-1}$ generates a contraction semigroup on $H$, implying $A=A_{\omega}+\omega I$ generates a quasicontraction semigroup with rate $\omega$ on $H$.
Now, for necessity, suppose $A$ generates a quasicontraction semigroup with rate $\omega$. Then $A_{\omega}:=A-\omega I$ generates a contraction semigroup. Defining $\mathcal{A}_{\omega}:=A_{\omega}\circ\mathcal{T}=\mathcal{A}-\omega\mathcal{T}$, the Lumer-Phillips theorem then implies $\left\langle\mathbf{u},A_{\omega}\mathbf{u}\right\rangle_{H}\leq 0$ for all $\mathbf{u}\in D$ and $\lambda_{\omega}I-A_{\omega}:D\to H$ is surjective for all $\lambda_{\omega}>0$. By Lem. 8, it follows that $\left\langle\mathcal{T}\mathbf{v},\mathcal{A}_{\omega}\mathbf{v}\right\rangle_{H}\leq 0$ for all $\mathbf{v}\in H$, whence $\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{H}\leq\omega\left\langle\mathcal{T}\mathbf{v},\mathcal{T}\mathbf{v}\right\rangle_{H}$ for all $\mathbf{v}\in H$. In addition, since $\mathcal{T}:D\to H$ is bijective, $\lambda_{\omega}\mathcal{T}-\mathcal{A}_{\omega}:D\to H$ is surjective for all $\lambda_{\omega}>0$. By Lem. 9, it follows that
$\left\lVert{(\lambda_{\omega}\mathcal{T}^{*}-\mathcal{A}_{\omega}^{*})\mathbf{v}}\right\rVert_{H}\geq\epsilon\left\lVert{\mathbf{v}}\right\rVert_{H},$
for all $\lambda_{\omega}>0$. Setting $\lambda=\lambda_{\omega}+\omega>\omega$, the stated conditions hold. ∎
Theorem 10 provides necessary and sufficient conditions for $\{\mathcal{T},\mathcal{A}\}$ to generate a quasicontraction semigroup on $H$, with given rate $\omega$. In the following section, we show how these conditions can be formulated as linear operator inequalities, to verify generation of an arbitrary $C_{0}$-semigroup on $L_{2}^{n}[a,b]$.
V A PIE-Based Generation Theorem for Arbitrary $C_{0}$-Semigroups
Building on Thm. 10, we now derive sufficient conditions under which a PIE-compatible PDE generates a $C_{0}$-semigroup on $L_{2}^{n}[a,b]$, satisfying an exponential bound as defined below.
Definition 11
For a Hilbert space $H$, a $C_{0}$-semigroup $(S(t))_{t\geq 0}$ on $H$ has growth rate $\omega\in\mathbb{R}$ and gain $M\geq 1$ if
$\left\lVert{S(t)}\right\rVert_{\mathcal{L}(H)}\leq Me^{\omega t},\qquad\forall t\geq 0.$
It is well-known that any $C_{0}$-semigroup satisfies a bound of this form for some $\omega\in\mathbb{R}$ and $M\geq 1$ [2]. However, while Thm. 10 allows arbitrary growth rates $\omega$, it restricts the resulting semigroup to have gain $M=1$. In the following subsection, therefore, we establish conditions for generation of a $C_{0}$-semigroup with arbitrary gain $M$ on $L_{2}^{n}$, by testing generation of a quasicontraction semigroup in the weighted inner product space, $H=(L_{2}^{n},\left\langle,\right\rangle_{\mathcal{P}})$. We then show how these conditions can be verified via convex optimization by parameterizing $\mathcal{P}$ as a partial integral operator.
V-A Generation of $C_{0}$-Semigroups on $L_{2}$
Thm. 10 provides necessary and sufficient conditions for $\{\mathcal{T},\mathcal{A}\}$ to generate a quasicontraction semigroup on a Hilbert space $H$. Applying this result directly to $H=L_{2}^{n}$, then, this only certifies generation of a $C_{0}$-semigroup with gain $M=1$. To capture more general semigroups, we instead consider the Hilbert space $H=(L_{2}^{n},\left\langle\cdot,\cdot\right\rangle_{\mathcal{P}})$ with weighted inner product $\left\langle\mathbf{u},\mathbf{v}\right\rangle_{\mathcal{P}}:=\left\langle\mathbf{u},\mathcal{P}\mathbf{v}\right\rangle_{L_{2}}$, for $\mathcal{P}:L_{2}^{n}\to L_{2}^{n}$. Here, choosing $\mathcal{P}$ to be coercive and bounded, the weighted inner product is equivalent to the standard $L_{2}$ inner product, allowing generation of a $C_{0}$-semigroup on $L_{2}^{n}$ to be verified by testing generation of a quasicontraction semigroup on $(L_{2}^{n},\left\langle\cdot,\cdot\right\rangle_{\mathcal{P}})$, yielding the following result.
Theorem 12
For given $n\in\mathbb{N}$ and $a,b]\subseteq\mathbb{R}$, let $\mathcal{T}:L_{2}^{n}[a,b]\to D$ and $\mathcal{A}:\in\mathcal{L}(L_{2}^{n}[a,b])$ be bounded linear operators, with $\mathcal{T}$ bijective onto $D$. Define $A:=\mathcal{A}\circ\mathcal{T}^{-1}:D\to L_{2}^{n}[a,b]$. For any $\omega\in\mathbb{R}$, if the following conditions are satisfied,
  1. 1.
    there exist $\epsilon_{1}>0$ and $\mathcal{P}\in\mathcal{L}(L_{2}^{n}[a,b])$ such that $\mathcal{P}\succeq\epsilon_{1}^{2}I$ and
    $\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{\mathcal{P}}\leq\omega\left\langle\mathcal{T}\mathbf{v},\mathcal{T}\mathbf{v}\right\rangle_{\mathcal{P}},\qquad\forall\mathbf{v}\in L_{2}^{n}[a,b];$
  2. 2.
    there exist $\epsilon_{2}>0$ and $\lambda>\omega$ such that
    $\left\lVert{(\lambda\mathcal{T}^{*}-\mathcal{A}^{*})\mathbf{v}}\right\rVert_{L_{2}}\geq\epsilon_{2}\left\lVert{\mathbf{v}}\right\rVert_{L_{2}},\qquad\forall\mathbf{v}\in L_{2}^{n}[a,b];$
then $A$ generates a $C_{0}$-semigroup on $L_{2}$ with growth rate $\omega$ and gain $M:=\frac{\sqrt{\left\lVert{\mathcal{P}}\right\rVert_{\mathcal{L}(L_{2})}}}{\epsilon_{1}}$.
Proof:
Suppose both conditions hold. Then,
$\epsilon_{1}^{2}\left\lVert{\mathbf{v}}\right\rVert_{L_{2}}^{2}\leq\left\lVert{\mathbf{v}}\right\rVert_{\mathcal{P}}^{2}\leq\left\lVert{\mathcal{P}}\right\rVert_{\mathcal{L}(L_{2})}\left\lVert{\mathbf{v}}\right\rVert_{L_{2}}^{2},\quad\forall\mathbf{v}\in L_{2}^{n}.$
Defining $\epsilon:=\frac{\epsilon_{2}}{\sqrt{\left\lVert{\mathcal{P}}\right\rVert_{\mathcal{L}(L_{2})}}}$, it follows that for all $\mathbf{v}\in L_{2}^{n}$,
$\left\lVert{(\lambda\mathcal{T}^{*}-\mathcal{A}^{*})\mathbf{v}}\right\rVert_{L_{2}}^{2}\geq\epsilon_{2}^{2}\left\lVert{\mathbf{v}}\right\rVert_{L_{2}}^{2}\geq\epsilon^{2}\left\lVert{\mathbf{v}}\right\rVert_{\mathcal{P}}^{2}.$
By Thm. 10, $A$ generates a semigroup on $H:=(L_{2},\left\langle\cdot,\cdot\right\rangle_{\mathcal{P}})$ with growth rate $\omega$. This contraction semigroup is also a $C_{0}$-semigroup on $(L_{2},\left\langle\cdot,\cdot\right\rangle_{L_{2}})$, where for all $\mathbf{x}\in D$,
$\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}\leq\frac{1}{\epsilon_{1}}\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{\mathcal{P}}\leq\frac{1}{\epsilon_{1}}e^{\omega t}\left\lVert{\mathbf{x}}\right\rVert_{\mathcal{P}}\leq Me^{\omega t}\left\lVert{\mathbf{x}}\right\rVert_{L_{2}}.$
Theorem 12 provides sufficient conditions for $\{\mathcal{T},\mathcal{A}\}$ to generate a $C_{0}$-semigroup with arbitrary gain $M$ and rate $\omega$ on $L_{2}^{n}$. The condition $\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{\mathcal{P}}\leq\omega\left\langle\mathcal{T}\mathbf{v},\mathcal{T}\mathbf{v}\right\rangle_{\mathcal{P}}$ herein is closely related to Lyapunov conditions for exponential stability in the PIE framework, differing only in that well-posedness does not require $\omega\leq 0$. Optimization-based methods for testing this condition have already been proposed for stability analysis in the PIE representation [14]. In the following subsection, we show how the well-posedness conditions may be similarly tested using convex optimization methods.
V-B An LPI for Well-Posedness of PDEs
Having established a test for well-posedness of PIEs in Thm. 12, we now show how this result can be applied to PDEs of the form in (4), on PIE-compatible domains $D$ as in Defn. 2. For such systems, we can define an associated PIE by $\mathcal{T}$ as in Cor. 3 and $\mathcal{A}:=A\circ\mathcal{T}$ as in Eqn. (4). It has been shown that both $\mathcal{T}$ and $\mathcal{A}$ belong to the class of Partial Integral (PI) operators [14], taking the form
$(\mathcal{P}\mathbf{v})(s)$ $:=\boldsymbol{P}_{0}(s)\mathbf{v}(s)+\int_{0}^{s}\boldsymbol{P}_{1}(s,\theta)\mathbf{v}(\theta)$
$\qquad+\int_{s}^{1}\boldsymbol{P}_{2}(s,\theta)\mathbf{v}(\theta)d\theta,$
and parameterized by polynomials $\boldsymbol{P}_{0},\boldsymbol{P}_{1},\boldsymbol{P}_{2}$. Such PI operators form a class of bounded, linear operators, that it closed under adjoint and composition, with explicit formulae for these operations given in [14]. In addition, PI operator variables can be parameterized using a finite monomial basis, and positivity and equality constraints can be enforced via semidefinite programming. Using these properties, the following corollary shows how well-posedness in the PIE representation can be verified by solving a convex optimization problem on PI operator variables.
Corollary 13
For given $n,d\in\mathbb{N}$ and $B,C\in\mathbb{R}^{2n\times 4n}$, let the associated domain $D\subseteq W_{2}^{d,n}[a,b]$ as in Defn. 2 be PIE admissible. For any $\boldsymbol{A}_{i}\in\mathbb{R}^{n\times n}[s]$, define $A:D\to L_{2}^{n}[a,b]$ as in (4). Define $\mathcal{T}\in\mathcal{L}(L_{2})$ as in Cor. 3, and let $\mathcal{A}:=A\circ\mathcal{T}\in\mathcal{L}(L_{2})$. For any $\epsilon_{1},\epsilon_{2}>0$, if there exists a PI operator $\mathcal{P}$ and scalar $\omega\in\mathbb{R}$ solving the optimization program
$\mathcal{P}\succeq\epsilon_{1}^{2},$ (6)
$\mathcal{T}^{*}\mathcal{P}\mathcal{A}+\mathcal{A}^{*}\mathcal{P}\mathcal{T}\preceq 2\omega\mathcal{T}^{*}\mathcal{P}\mathcal{T},$
$([\omega+1]\mathcal{T}-\mathcal{A})([\omega+1]\mathcal{T}^{*}-\mathcal{A}^{*})\succeq\epsilon_{2}^{2}I,$
then $A$ is the infinitesimal generator of a $C_{0}$-semigroup on $L_{2}^{n}[a,b]$ with growth rate $\omega$ and gain $M:=\frac{\sqrt{\left\lVert{\mathcal{P}}\right\rVert_{\mathcal{L}(L_{2})}}}{\epsilon_{1}}$.
Proof:
Suppose $(\mathcal{P},\omega)$ solves the optimization program. Then, for all $\mathbf{v}\in L_{2}^{n}[a,b]$,
$\left\langle\mathcal{T}\mathbf{v},\mathcal{A}\mathbf{v}\right\rangle_{\mathcal{P}}=\frac{1}{2}\left\langle\mathbf{v},[\mathcal{T}^{*}\mathcal{P}\mathcal{A}+\mathcal{A}^{*}\mathcal{P}\mathcal{T}]\mathbf{v}\right\rangle_{L_{2}}\leq\omega\left\langle\mathcal{T}\mathbf{v},\mathcal{T}\mathbf{v}\right\rangle_{\mathcal{P}}.$
In addition, defining $\lambda:=\omega+1>\omega$, we obtain
$\left\lVert{(\lambda\mathcal{T}^{*}-\mathcal{A}^{*})\mathbf{v}}\right\rVert_{L_{2}}^{2}\geq\epsilon_{2}^{2}\left\lVert{\mathbf{v}}\right\rVert_{L_{2}}^{2}.$
By Thm. 12, it follows that $A$ is the infinitesimal generator of a $C_{0}$-semigroup on $L_{2}^{n}[a,b]$ with growth rate $\omega$ and gain $M:=\frac{\sqrt{\left\lVert{\mathcal{P}}\right\rVert_{\mathcal{L}(L_{2})}}}{\epsilon_{1}}$. ∎
By Cor. 13, well-posedness of a PIE-compatible PDE can be tested by solving the optimization program in (6). For each fixed $\omega$, this is a feasibility problem that is linear in the decision variable $\mathcal{P}$ and defined by operator inequalities on PI operators. We refer to such a program as a Linear PI Inequality (LPI). Using the PIETOOLS software suite [16], this LPI can be converted to a semidefinite program and solved with an SDP solver such as Mosek [20]. Using bijection, a smallest value of $\omega$ may then be established for which the program is feasible, thereby certifying well-posedness together with a least upper bound on the exponential growth rate. In the next section, we apply this optimization program to several PDE examples and obtain tight upper bounds on the exponential growth rate of solutions.
Well-Posedness in a different Hilbert-Space
Although Cor. 13 allows well-posedness to be verified of a broad class of PDEs, it only tests for generation of a $C_{0}$-semigroup on the Hilbert space $L_{2}^{n}[a,b]$, with standard inner product. In particular, the constraint $\mathcal{T}^{*}\mathcal{P}\mathcal{A}+\mathcal{A}^{*}\mathcal{P}\mathcal{T}\preceq 2\omega\mathcal{T}^{*}\mathcal{P}\mathcal{T}$ for $\mathcal{P}\succeq\epsilon^{2}I$ ensures that solutions satisfy $\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}\leq Me^{\omega t}\left\lVert{\mathbf{x}}\right\rVert_{L_{2}}$ for some $M\geq 1$, as is necessary for $e^{tA}$ to be a $C_{0}$-semigroup on $L_{2}^{n}$, However, many interesting PDEs are not well-posed on $L_{2}^{n}$ with the standard inner product (see, e.g., Subsec. VI-C 2)). In such cases, solutions may instead satisfy a weaker bound of the form $\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}\leq e^{\omega t}\left\lVert{\partial_{s}^{d}\mathbf{x}}\right\rVert_{L_{2}}$, which can be tested using non-coercive operators $\mathcal{P}\succeq\epsilon\mathcal{T}$ [21]. An open question is whether a similar modification of the conditions of Thm. 12 may be used to test well-posedness with such an alternative bound on solutions, verifying generation of a $C_{0}$-semigroup on a different Hilbert space.
VI Examples
Given the well-posedness test for linear PDEs in Cor. 13, we now verify the efficacy of this test for verifying well-posedness of several classical PDE examples, as well as a more complicated system. For each example, the optimization program in (6) is implemented in PIETOOLS [16], with $\epsilon_{1}=\epsilon_{2}=10^{-1}$, and the resulting semidefinite program is solved using Mosek [20]. For each of the classical PDE examples, an ill-posed version is also formulated by modifying either the definition of $A$ or of its domain, and we verify that the LPI in (6) becomes infeasible. Analytic proofs of well-posedness for the examples are provided in Appx. A.
VI-A Transport Equation
VI-A1 Well-Posed Example
As a first example, consider the transport equation
$u_{t}(t,s)=-u_{s}(t,s),\qquad u(t,0)=0.$ (7)
We formulate an associated Cauchy problem as in (3) on $H:=L_{2}[0,1]$, by defining $A:D\to L_{2}[0,1]$ as
$A:=-\partial_{s},\qquad D:=\bigl\{\mathbf{u}\in W_{2}^{1}[0,1]\mid\mathbf{u}(0)=0\bigr\}.$
Then $A$ is the infinitesimal generator of a $C_{0}$-semigroup on $H$, with exponential growth rate $\omega$ and gain $M=e^{-\omega}$ for any $\omega\leq 0$. To verify well-posedness using Cor. 13, define PI operators $(\mathcal{T}\mathbf{v})(s):=\int_{0}^{s}\mathbf{v}(\theta)d\theta$ and $\mathcal{A}:=-I$. Then $\{\mathcal{T},\mathcal{A}\}$ define the equivalent PIE representation, with $\mathcal{T}^{-1}=\partial_{s}$ and $\mathcal{A}=A\circ\mathcal{T}$. Solving the optimization program in (6) with $\mathcal{P}=I$, well-posedness of the PDE can be verified with $\omega=0$, certifying that $A$ generates a contraction semigroup. Furthermore, parameterizing a more general PI operator variable $\mathcal{P}$ using monomials of degree at most $2d$, a least upper bound on the exponential growth rate $\omega$ of the semigroups is established as in Table I.
$d$ 0 1 2 3 4 5 6 7 8
$\omega$ 0 -0.999 -2.366 -3.874 -4.834 -4.958 -5.067 -5.079 -5.100
TABLE I: Least upper bound on the exponential growth rate $\omega$ with which well-posedness of the transport equation in (7) can be verified by solving the LPI in Cor. 13, parameterizing $\mathcal{P}$ by monomials of degree at most $2d$.
VI-A2 Ill-Posed Example
Consider now the transport equation with boundary condition applied at the opposite boundary,
$u_{t}(t,s)=-u_{s}(t,s),\qquad u(t,1)=0.$
In this case the PDE is not well-posed, admitting no solution for nontrivial initial states. As such, defining the associated PIE representing, setting $(\mathcal{T}\mathbf{v})(s)=-\int_{s}^{1}\mathbf{v}(\theta)d\theta$ and $\mathcal{A}:=-I$, the LPI (6) is infeasible for all $\omega\in\mathbb{R}$.
VI-B Heat Equation
VI-B1 Well-Posed Example
Consider the reaction–diffusion equation
$u_{t}(t,s)$ $=u_{ss}(t,s)+ru(t,s),$ $t\geq 0,\ s\in[0,1],$
$u(t,0)$ $=u(t,1)=0,$
where $r\in\mathbb{R}$. Representing this PDE as in (3) with
$A:=\partial_{s}^{2}+rI,\quad D=\bigl\{\mathbf{u}\in W_{2}^{2}[0,1]\mid\mathbf{u}(0)=\mathbf{u}(1)=0\bigr\},$
AA generates a $C_{0}$-semigroup with growth rate $\omega=r-\pi^{2}$ and gain $M=1$ on $L_{2}[0,1]$, for all $r\in\mathbb{R}$. We can represent this system as a PIE by defining
$(\mathcal{T}\mathbf{v})(s)=\int_{0}^{s}\theta(s-1)\mathbf{v}(\theta)d\theta+\int_{s}^{1}s(\theta-1)\mathbf{v}(\theta)d\theta,$ (8)
and $\mathcal{A}:=A\circ\mathcal{T}=I+r\mathcal{T}$. Solving the LPI in (6), parameterizing $\mathcal{P}$ by polynomials of degree at most $6$, well-posedness can be verified for any $r\in\{0,1,\dots,20\}$, with least upper bound $\omega=\pi^{2}-r-10^{-5}$ on the exponential growth rate.
VI-B2 Ill-Posed Example
Consider the reverse-time heat equation
$u_{t}(t,s)$ $=-u_{ss}(t,s),$ $t\geq 0,\ s\in[0,1],$
$u(t,0)$ $=u(t,1)=0.$
Defining $A=-\partial_{s}^{2}$ on the same domain $D$ as before, the resulting Cauchy problem is not well-posed. Consequently, defining the associated PIE representation by $\mathcal{T}$ as in (8) and $\mathcal{A}=-I$, the LPI (6) is infeasible for all $\omega\in\mathbb{R}$.
VI-C Wave Equation
VI-C1 Well-Posed Example
Now, consider the wave equation
$u_{tt}(t,s)$ $=u_{ss}(t,s),$ $t\geq 0,\ s\in[0,1],$ (9)
$u_{t}(t,0)$ $=u_{s}(t,1)=0.$ (10)
Defining $\mathbf{u}(t)=(u_{t}(t),u_{s}(t))$, this PDE can be written as in (3) with
$A=\begin{bmatrix}0&\partial_{s}\\ \partial_{s}&0\end{bmatrix},\quad D=\Bigl\{\mathbf{u}\in W_{2}^{1,2}[0,1]\,\Bigl|\ \begin{matrix}\mathbf{u}_{1}(0)=0\\ \mathbf{u}_{2}(1)=0\end{matrix}\,\Bigr\}.$
Then, $A$ generates a contraction semigroup on $L_{2}^{2}[0,1]$. For this system, the equivalent PIE is defined by
$(\mathcal{T}\mathbf{v})(s)=\int_{0}^{s}\begin{bmatrix}1&0\\ 0&0\end{bmatrix}\mathbf{v}(\theta)d\theta-\int_{s}^{1}\begin{bmatrix}0&0\\ 0&1\end{bmatrix}\mathbf{v}(\theta)d\theta,$
with $\mathcal{T}^{-1}=\partial_{s}$, and $\mathcal{A}=A\circ\mathcal{T}={{\scriptsize\left[\!\!\begin{array}[]{cc}0&1\\ 1&0\end{array}\!\!\right]}}$. The optimization program (6) is feasible with $\omega=0$ and $\mathcal{P}=I$, certifying that $A$ generates a contraction semigroup.
VI-C2 Ill-Posed Example
For the same wave equation in (9), suppose we instead define the state $\mathbf{u}(t)=(u(t),u_{t}(t))$ with
$A=\begin{bmatrix}0&I\\ \partial_{s}^{2}&0\end{bmatrix},\quad D=\Bigl\{\mathbf{u}\in W_{2}^{2,2}[0,1]\,\Bigl|\ \begin{matrix}\mathbf{u}(0)=0\\ \mathbf{u}_{s}(1)=0\end{matrix}\,\Bigr\}.$
Then, the resulting Cauchy problem is not well-posed. Computing $\mathcal{T}=(\partial_{s}^{2})^{-1}$ and $\mathcal{A}=A\circ\mathcal{T}$ yields a PIE representation whose operator $\mathcal{Q}:=(\lambda\mathcal{T}-\mathcal{A})(\lambda\mathcal{T}-\mathcal{A})^{*}$ cannot satisfy $\mathcal{Q}\succeq\epsilon I$ for any $\epsilon>0$ or $\lambda\in\mathbb{R}$. Consequently, the LPI (6) is infeasible for all $\omega\in\mathbb{R}$.
VI-D Coupled PDEs
To illustrate the application of Cor. 13 for verifying well-posedness of less classically well-studied PDEs, consider finally the following system of a coupled PDEs
$\partial_{t}u_{1}(t,s)$ $=-\partial_{s}u_{1}(t,s)+s\partial_{t}u_{3}(t,1),$
$\partial_{t}u_{2}(t,s)$ $=\partial_{s}^{2}u_{2}(t,s)+s(2-s)u_{1}(t,1)+\partial_{s}u_{3}(t,0),$
$\partial_{t}^{2}u_{3}(t,s)$ $=\partial_{s}^{2}u_{3}(t,s)-\partial_{s}u_{2}(t,s),$
$u_{1}(t,0)$ $=u_{2}(t,1)=\partial_{t}u_{3}(t,0),\qquad u_{2}(t,0)=0,$
$\partial_{s}u_{2}(t,1)$ $=\partial_{s}u_{3}(t,0),\hskip 56.9055pt\partial_{s}u_{3}(t,1)=0.$
Introducing the state $\mathbf{u}(t)=(u_{1}(t),u_{2}(t),\partial_{s}u_{3}(t),\partial_{t}u_{3}(t))$, this system can be formulated as in (3) using a suitable operator $A:D\to L_{2}^{4}[0,1]$ on the domain
$D:=\biggl\{\mathbf{u}\in{\small\left[\!\begin{array}[]{l}W_{2}^{1}[0,1]\\ W_{2}^{2}[0,1]\\ W_{2}^{1,2}[0,1]\end{array}\!\right]}\,\Bigl|\ {\small\begin{array}[]{l}\mathbf{u}_{1}(0)=\mathbf{u}_{2}(1)=\mathbf{u}_{3}(0)\\ \mathbf{u}_{2}(0)=\mathbf{u}_{4}(1)=0\\ \partial_{s}\mathbf{u}_{2}(1)=\mathbf{u}_{4}(0)\end{array}}\biggr\}$
An explicit expression for $A$ and the operators $\{\mathcal{T},\mathcal{A}\}$ defining the associated PIE representation are omitted for brevity. Solving the LPI in (6), we can verify that this PDE is well-posed, generating a $C_{0}$-semigroup with growth rate $\omega=0$.
VII Conclusion
A PIE-based framework for verifying well-posedness of partial differential equations was developed. By reformulating the Lumer–Phillips generation theorem in terms of Partial Integral (PI) operators, necessary and sufficient conditions were established for a PIE to generate a quasicontraction semigroup. These conditions were relaxed to determine sufficient conditions for generation of a more general $C_{0}$-semigroup, with upper bound on the exponential growth rate of solutions. Using these results, a convex optimization program on PI operator variables was developed for testing well-posedness of any PDE admitting an equivalent PIE representation. The resulting conditions can be implemented using PIETOOLS and solved using semidefinite programming. Numerical examples demonstrated that the method can verify well-posedness and accurately estimate growth rates for several 1D PDE systems.
Appendix A Well-Posedness of Classical PDE Examples
In this appendix, we consider the classical PDE examples from Sec. VI, and verify well-posedness properties. In each case, we will demonstrate well-posedness by explicitly defining the $C_{0}$-semigroup generated by the PDE, or proving that such a $C_{0}$-semigroup does not exist.
A-A Transport Equation
Consider first the operator
$A:=-\partial_{s}:D\to L_{2}[0,1],$
$D:=\{\mathbf{u}\in W_{2}^{1}[0,1]\mid\mathbf{u}(0)=0\},$
defining a transport equation with Dirichlet boundary condition at the lower boundary. A solution to the transport equation can be readily found using the method of characteristics, noting that $u_{t}(t,s)=-u_{s}(t,s)$ and $u(0,s)=\mathbf{x}(s)$ implies $u(t,s)=\mathbf{x}(s-t)$ for all $t\leq s$. Imposing the boundary condition, $u(t,0)$, it follows that $A$ is the infinitesimal generator of the following $C_{0}$-semigroup.
$(e^{tA}\mathbf{x})(s):=\begin{cases}\mathbf{x}(s-t),&t\leq s,\\ 0,&t>s.\end{cases}$
Here, $\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}\leq\left\lVert{e^{t^{\prime}A}\mathbf{x}}\right\rVert_{L_{2}}$ for all $t\geq t^{\prime}\geq 0$ and $\mathbf{x}\in D$, whence $(e^{tA})_{t\geq 0}$ is a contraction semigroup. Furthermore since $\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}=0$ for all $t\geq 1$ and $\mathbf{x}\in D$, it follows that $\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}\leq e^{\omega[t-1]}\left\lVert{\mathbf{x}}\right\rVert_{L_{2}}$ for any $\omega\in\mathbb{R}$ all $t\geq 1$ and $\mathbf{x}\in D$. Thus, the $C_{0}$-semigroup generated by $A$ has exponential growth rate $\omega$ for any $\omega\in\mathbb{R}$, with gain $M=e^{\omega}$.
However, consider now the following operator,
$A:=-\partial_{s}:D\to L_{2}[0,1],$
$D:=\{\mathbf{u}\in W_{2}^{1}[0,1]\mid\mathbf{u}(1)=0\},$
imposing a Dirichlet boundary condition on the upper boundary. In this case, the semigroup generated by $A$ would still need to satisfy $(e^{tA}\mathbf{x})(s)=\mathbf{x}(t-s)$ for $t\leq s$. However, for any nontrivial $\mathbf{x}\in D$, we may not have $\mathbf{x}(t-1)=0$ for arbitrary $t\geq 0$, implying that $e^{tA}\mathbf{x}\notin D$ for all $t\geq 0$. Therefore, $A$ cannot be the infinitesimal generator of a $C_{0}$-semigroup, implying the associated transport equation is ill-posed.
A-B Heat Equation
Next, consider the operator
$A:=\partial_{s}^{2}:D\to L_{2}[0,1],$
$D:=\{\mathbf{u}\in W_{2}^{2}[0,1]\mid\mathbf{u}(0)=\mathbf{u}(1)=0\},$
defining the heat equation with Dirichlet boundary conditions. The solution to the heat equation can be readily established using separation of variables, finding that $A$ generates the $C_{0}$-semigroup
$(e^{At}\mathbf{x})(s)=\sum_{n=1}^{\infty}2\left\langle\mathbf{x},\boldsymbol{\phi}_{n}\right\rangle_{L_{2}}e^{-n^{2}\pi^{2}t}\boldsymbol{\phi}_{n}(s),$
where $\boldsymbol{\phi}_{n}(s):=\sin(n\pi s)$ for $n\in\mathbb{N}$. However, consider now the reverse-time heat equation, defined by
$A:=-\partial_{s}^{2}:D\to L_{2}[0,1],$
$D:=\{\mathbf{u}\in W_{2}^{2}[0,1]\mid\mathbf{u}(0)=\mathbf{u}(1)=0\}.$
It is clear that if this operator generates a $C_{0}$-semigroup, then this semigroup would have to be given by
$(e^{At}\mathbf{x})(s)=\sum_{n=1}^{\infty}2\left\langle\mathbf{x},\boldsymbol{\phi}_{n}\right\rangle_{L_{2}}e^{n^{2}\pi^{2}t}\boldsymbol{\phi}_{n}(s).$
However, this semigroup is not continuous, since for any $M>0$ and $t>0$, we can establish $n\in\mathbb{N}$ sufficiently large such that $e^{n\pi^{2}t}>2M$, whence $\left\lVert{e^{At}\boldsymbol{\phi}_{n}}\right\rVert_{L_{2}}>M$. Thus, the operator $A$ does not generate a $C_{0}$-semigroup, implying the associated reverse-time heat equation is not well-posed.
A-C Wave Equation
Consider the wave equation with mixed boundary conditions,
$u_{tt}(t,s)$ $=u_{ss}(t,s),\qquad t\geq 0,~s\in[0,1],$
$u(t,0)$ $=u_{s}(t,1)=0$
For this equation, solutions are given by
$u(t,s)$ $:=\sum_{n=0}^{\infty}\bigl(a_{n}\cos(\mu_{n}t)+b_{n}\sin(\mu_{n}t)\bigr)\boldsymbol{\phi}_{n}(s)$
where $\mu_{n}:=[n+\frac{1}{2}]\pi$ and $\boldsymbol{\phi}_{n}(s):=\sin(\mu_{n}s)$ for $n\in\mathbb{N}$, and where
$a_{n}$ $:=2\left\langle\boldsymbol{\phi}_{n},u(0,\cdot)\right\rangle_{L_{2}}=2\int_{0}^{1}\sin(\mu_{n}s)u(0,s)\,ds,$
$b_{n}$ $:=2\left\langle\boldsymbol{\phi}_{n},u_{t}(0,\cdot)\right\rangle_{L_{2}}=2\int_{0}^{1}\sin(\mu_{n}s)u_{t}(0,s)\,ds.$
It follows that
$u_{t}(t,s)$ $:=\sum_{n=0}^{\infty}\mu_{n}\bigl(-a_{n}\sin(\mu_{n}t)+b_{n}\cos(\mu_{n}t)\bigr)\boldsymbol{\phi}_{n}(s),$
$u_{s}(t,s)$ $:=\sum_{n=0}^{\infty}\mu_{n}\bigl(a_{n}\cos(\mu_{n}t)+b_{n}\sin(\mu_{n}t)\bigr)\boldsymbol{\phi}_{n}^{\prime}(s),$
where $\boldsymbol{\phi}_{n}^{\prime}(s):=\cos(\mu_{n}(s))$.
Now, to formulate the wave equation as an abstract Cauchy problem as in (3), we may introduce the state $\mathbf{u}(t)=\begin{bmatrix}u_{t}(t)\\ u_{s}(t)\end{bmatrix}$, defining
$A:=\begin{bmatrix}0&\partial_{s}\\ \partial_{s}&0\end{bmatrix}:D\to L_{2}^{2}[0,1],$
$D:=\{(\mathbf{u}_{1},\mathbf{u}_{2})\in W_{2}^{1,2}[0,1]\mid\mathbf{u}_{1}(0)=\mathbf{u}_{2}(1)=0\}.$
Then, the $C_{0}$-semigroup generated by $A$ must satisfy
$(e^{At}\mathbf{x})(s):=\sum_{n=0}^{\infty}\mu_{n}\begin{bmatrix}\bigl(-a_{n}\sin(\mu_{n}t)+b_{n}\cos(\mu_{n}t)\bigr)\boldsymbol{\phi}_{n}(s)\\ \bigl(a_{n}\cos(\mu_{n}t)+b_{n}\sin(\mu_{n}t)\bigr)\boldsymbol{\phi}_{n}^{\prime}(s)\end{bmatrix}$.
where now $a_{n}:=2\left\langle\boldsymbol{\phi}_{n},\mathbf{x}_{1}\right\rangle_{L_{2}}$ and $b_{n}:=2\left\langle\boldsymbol{\phi}_{n},\mathbf{x}_{2}\right\rangle_{L_{2}}$ for $\mathbf{x}=\begin{bmatrix}\mathbf{x}_{1}\\ \mathbf{x}_{2}\end{bmatrix}\in D$. By orthogonality of the basis functions, it follows that
$\left\lVert{e^{At}\mathbf{x}}\right\rVert_{L_{2}}^{2}$ $=\sum_{n=0}^{\infty}\frac{\mu_{n}^{2}}{2}\Bigl[\bigl(-a_{n}\sin(\mu_{n}t)+b_{n}\cos(\mu_{n}t)\bigr)^{2}$
$\qquad+\bigl(a_{n}\cos(\mu_{n}t)+b_{n}\sin(\mu_{n}t)\bigr)^{2}\Bigr]$
$=\sum_{n=0}^{\infty}\frac{\mu_{n}^{2}}{2}\bigl[a_{n}^{2}+b_{n}^{2}\bigr]=\left\lVert{\mathbf{x}}\right\rVert_{L_{2}}^{2}.$
Hence, $e^{At}$ is a contraction semigroup.
Alternatively, the wave equation can also be represented as in (3) by introducing $\mathbf{u}(t)=\begin{bmatrix}u(t)\\ u_{t}(t)\end{bmatrix}$ and defining
$A:=\begin{bmatrix}0&I\\ \partial_{s}^{2}&0\end{bmatrix}:D\to L_{2}^{2}[0,1],$
$D:=\biggl\{\mathbf{u}\in W_{2}^{2,2}[0,1]\,\Bigl|\ \begin{matrix}\mathbf{u}_{1}(0)=\partial_{s}\mathbf{u}_{1}(1)=0\\ \mathbf{u}_{2}(0)=\partial_{s}\mathbf{u}_{2}(1)=0\end{matrix}\biggr\}.$
Then, if $A$ is the generator of a $C_{0}$-semigroup, this semigroup must satisfy
$(e^{tA}\mathbf{x})(s):=\sum_{n=0}^{\infty}2\begin{bmatrix}a_{n}\cos(\mu_{n}t)+b_{n}\sin(\mu_{n}t)\\ \mu_{n}\bigl(a_{n}\sin(\mu_{n}t)+b_{n}\cos(\mu_{n}t)\bigr)\end{bmatrix}\boldsymbol{\phi}_{n}(s),$
where again $a_{n}:=2\left\langle\boldsymbol{\phi}_{n},\mathbf{x}_{1}\right\rangle_{L_{2}}$ and $b_{n}:=2\left\langle\boldsymbol{\phi}_{n},\mathbf{x}_{2}\right\rangle_{L_{2}}$ for $\mathbf{x}=\begin{bmatrix}\mathbf{x}_{1}\\ \mathbf{x}_{2}\end{bmatrix}\in D$. By orthogonality of the basis functions, it follows that
$\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}^{2}$ $=\sum_{n=0}^{\infty}\frac{1}{2}\Bigl[\bigl(a_{n}\cos(\mu_{n}t)+b_{n}\sin(\mu_{n}t)\bigr)^{2}$
$\hskip 42.67912pt+\mu_{n}^{2}\bigl(-a_{n}\sin(\mu_{n}t)+b_{n}\cos(\mu_{n}t)\bigr)^{2}\Bigr].$
However, consider now $\mathbf{x}(s)=\begin{bmatrix}\sin(\mu_{m}s)\\ 0\end{bmatrix}$ for some $m\in\mathbb{N}$, so that $b_{n}:=0$ for all $n\in\mathbb{N}$, and $a_{n}=0$ for all $n\in\mathbb{N}\setminus\{m\}$, with $a_{m}=1$. Then
$\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}^{2}$ $=\frac{1}{2}\Bigl[\cos^{2}(\mu_{m}t)+\mu_{m}^{2}\sin^{2}(\mu_{m}t)\Bigr],$
whence $\left\lVert{\mathbf{x}}\right\rVert_{L_{2}}^{2}=1$ and $\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}^{2}|_{t=1}=\frac{1}{2}\mu_{m}^{2}$. It follows that, for any $M\geq 1$ and $\omega\in\mathbb{R}$, we can find sufficiently large $m\in\mathbb{N}$ such that $\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}^{2}|_{t=1}=\frac{1}{2}\mu_{m}^{2}>Me^{\omega}\left\lVert{\mathbf{x}}\right\rVert_{L_{2}}^{2}$. Since the $C_{0}$-semigroup generated by $A$ must satisfy $\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{H}\leq Me^{\omega t}\left\lVert{\mathbf{x}}\right\rVert_{H}$ for some $\omega\in\mathbb{R}$ and $M\geq 1$, it follows that the considered $A:D\to H$ cannot be the infinitesimal generator of a $C_{0}$-semigroup, wherefore this formulation of the wave equation cannot be well-posed.